OpenStudy (anonymous):

Integration! \[\int \sin^2 x dx\]

OpenStudy (anonymous):

use substituion to solve this one

OpenStudy (anonymous):

I actually think I got this one, but I also think my method is sort of crazy. \[u = \sin x\]\[du = \cos x\]\[dv = \sin x\]\[v = -\cos x\] \[uv - \int v du\]\[-\sin x\cos x + \int \cos^2 x dx\] \[\int \sin^2 x dx = -\sin x\cos x + \int 1 dx - \int \sin^2 x dx\]\[2\int \sin^2 x dx = -\sin x\cos x + \int 1 dx\] So I get \[\frac{x - \sin x\cos x}{2} + C \] Which is apparently the right answer. What is the "correct" way to do it?

OpenStudy (johnweldon1993):

Yeah that is definitely 1 way to do it...another would be to use the fact that \[\large \sin^2xdx = \frac{1}{2}(1 - \cos(2x))\] Integrate term by term \[\int\limits_{}^{}\frac{1}{2}dx - \int\limits_{}^{}\frac{1}{2}\cos(2x)dx\] \[\frac{1}{2}\int\limits_{}^{}dx - \frac{1}{2}\int\limits_{}^{}\cos(2x)dx\] \[\frac{x}{2} - \frac{1}{2}\int\limits_{}^{}\cos(2x)dx\] let u = 2x du = 2dx dx = du/2 \[\frac{x}{2} - \frac{1}{2}\int\limits_{}^{}\cos(u)\frac{du}{2}\] \[\frac{x}{2} - \frac{1}{4}\int\limits_{}^{}\cos(u)du\] \[\frac{x}{2} - \frac{1}{4}\sin (u)\] replace u = 2x \[\frac{x}{2} - \frac{1}{4}\sin (2x)\] we know \[\large \sin(2x) = 2\sin(x)\cos(x)\] so \[\frac{x}{2} - \frac{1}{2}\sin (x)cos(x)\] \[\frac{x - sin(x)cox(x)}{2} + C\]

OpenStudy (johnweldon1993):

*typo in that last one* \[\large \frac{x - \sin(x)\cos(x)}{2} + C\]

OpenStudy (anonymous):

I'm not sure if that's actually any shorter! But it might be a little more "sane." Maybe. Thank you.

OpenStudy (johnweldon1993):

Yeah I know what you mean...it requires more memorization of identities.